OpenStudy (anonymous):

find the value(s) of guaranteed by the Mean Value Theorem for Integrals for the function over the given interval. F(x)=9/x^3 [1,3]

OpenStudy (anonymous):

\[f(c)=\frac{9}{c^3}=\frac{1}{3-1}\int_1^3\frac{9}{x^3}\,dx\] \[c^3=\frac{2}{\displaystyle\int_1^3\frac{dx}{x^3}}\]